OpenStudy (anonymous):

Infinite Geometric Series. 1- 1/3 + 1/9 - ... find the sum

OpenStudy (anonymous):

3/4

OpenStudy (anonymous):

sum to infinity is 1 / (1-r) = 1 / (1-- 1/3) = 3/4